OpenStudy (anonymous):

Allison wants to be able to view two 8 1/2 X 11in. pages at full size side by side on her computer screen. She needs to view the pages with the 11 in. sides vertical and the 8 1/2 sides beside each other horizontally. What is the smallest diagonal screen size she should look for? Estimate your answer to one decimal place.

OpenStudy (anonymous):

Use Pythagorean Theorem. 17^2 + 11^2 = c^2 289 + 121 = c^2 c^2 = 410 c = √410 Just put √410 in your calculator if you want the approximate answer.

OpenStudy (anonymous):

where did the 17 come from

OpenStudy (anonymous):

8.5 + 8.5 = 17

OpenStudy (anonymous):

why do we need to add 8.5 + 8.5

OpenStudy (anonymous):

You need to add them to get the full length. THese are two rectangles lined up next to each other, so you need to add them to get the full length.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):